OpenStudy (jennychan12):

Nils is emptying the dishwasher. He removes a 0.200 kg glass that has a temperature of 30.0°C. Into it he pours 0.100 kg of diet soda (mostly water) which comes out of the refrigerator with a temperature of 5.00°C. Assuming no external heat loss, what will be the final equilibrium temperature of the glass of diet soda (no ice was added)? I got the answer to this: it's 12.29C. Nil doesn't feel that his drink is cold enough, so he throws in an ice cube of -3C. What is the pass of the ice cube if his drink and glass are now cooled to 1C? L = 4185 J/kg .I got 0.009 kg, but the answer is 0.0019 kg

OpenStudy (jennychan12):

Am i doing something wrong?

OpenStudy (jennychan12):

whoops i typed something wrong. L(f) = 3.3*10^5 c(w) = 4186 c(g) = 840

OpenStudy (shubhamsrg):

am getting 12.16 as the first ans. can you please share your work ?

OpenStudy (jennychan12):

i know the first answer is correct. 0.2(840)(30-T) = 0.1(4185)(T-5) T = 12.29C

OpenStudy (shubhamsrg):

yep..thats the correct eqn,,maybe just a slight diff since it should be 4186 and not 4185,,that doesnt matter ofcorse,,its correct..

OpenStudy (jennychan12):

i just don't understand how to set up the second problem. (Nil doesn't feel that his drink is cold enough, so he throws in an ice cube of -3C. What is the pass of the ice cube if his drink and glass are now cooled to 1C?)

OpenStudy (shubhamsrg):

what do you mean pass ? is that mass ?

OpenStudy (jennychan12):

sorry. i was typing really fast... :( yesh,mass

OpenStudy (shubhamsrg):

heat lost =heat gained heat will be lost by the glass and the drink,and gained by ice, firstly it'll melt and then temp change will occur..

OpenStudy (shubhamsrg):

(0.1)(4186)(12.16 - 1) + (0.2)(840)(12.16 -1) = m(3.3 * 10^5) + m(4186)(1 - (-3))

OpenStudy (shubhamsrg):

try it..and please get back to me,,am also not too confident with my ans! :P

OpenStudy (jennychan12):

thanks for trying to help! :D

OpenStudy (shubhamsrg):

did i actually was able to help? i get the ans as 0.019 kgs

OpenStudy (jennychan12):

that's the correct answer :)

OpenStudy (shubhamsrg):

phew!! so,,did you understand ? afterall,,i understood this right now only! :P

OpenStudy (jennychan12):

haha pretty good! yeah i understood it :3